# Composable Finite-Size Security of High-Dimensional Quantum Key Distribution Protocols

### v1

@@ Abstract
Practical implementations of Quantum Key Distribution (QKD) extending beyond urban areas commonly use satellite links. However, the transmission of quantum states through the Earth’s atmosphere is highly susceptible to noise, restricting its application primarily to nighttime. High-dimensional (HD) QKD offers a promising solution to this limitation by employing high-dimensionally entangled quantum states. Although experimental platforms for HD QKD exist, previous security analyses were limited to the asymptotic regime and have either relied on impractical measurements or employed computationally demanding convex optimization tasks restricting the security analysis to low dimensions. In this work, we bridge this gap by presenting a composable finite-size security proof against both collective and coherent attacks for a general HD QKD protocol that utilizes only experimentally accessible measurements. In addition to the conventional, yet impractical ‘one-shot’ key rates, we also provide a practical variable-length security argument that yields significantly higher expected key rates. This approach is particularly crucial for rapidly changing and turbulent atmospheric conditions, as encountered for free-space and satellite-based QKD platforms.
@@ I Introduction
Quantum Key Distribution [1, 2] enables two remote parties, called Alice and Bob, to establish cryptographically secure keys, without relying on assumptions on the hardness of mathematical problems. Over recent decades, significant advancements in both experimental techniques and security proofs evolved QKD from an interesting scientific concept towards one of the most developed quantum technologies [3, 4, 5], being on the edge of practical use for both private and governmental applications. Implementations range from fiber links [6, 7, 8, 9, 10] for urban areas to free-space [11, 12] and satellite-based links [13, 14, 15, 16, 17] that span countries and even continents. Although the transmission of quantum signals through the atmosphere offers a more favorable loss profile compared to fiber-based systems, it also suffers from high noise sensitivity, which basically confines operation hours to nighttime.
@@
High-dimensional quantum states [18, 19, 20] have proven to enhance noise tolerance in entanglement distribution tasks [21], making, high-dimensional quantum key distribution (HD QKD) a promising candidate for improving noise robustness in cryptographic tasks. Nonetheless, the security analysis of practically implementable HD QKD protocols is so far rather uncharted. Existing works primarily focus on asymptotic security, employing either analytical methods [22, 23, 24, 25, 26] that leverage symmetry arguments or numerical approaches [27] involving convex optimization [28, 29, 30]. Analytical arguments often require $2$ or even $d+1$ mutually unbiased bases (MUB) measurements, which are impractical, particularly for large dimensions. Numerical methods avoid this but require solving very high-dimensional convex optimization problems which are computationally demanding and very memory-intensive, rendering them infeasible even for low to medium dimensions [27]. Recent work by Ref. [31] has closed this gap between experimentally available setups [32, 33, 34, 35, 36] and asymptotic security, by providing an efficient semi-analytic method for evaluating key rates for practical HD QKD setups. Yet, the security of high-dimensional QKD protocols in the finite-size regime remains an open question.
@@
In this study, we address this issue by presenting a full composable finite-size security proof for general high-dimensional QKD protocols. We begin by outlying the general HD QKD protocol in Section II. Section III details our statistical testing procedure and establishes composable finite-size security against i.i.d. collective attacks. We then apply the postselection technique [37, 38] to lift to security against coherent attacks. These secure key rates refer to the so-called fixed-length scenario, which is commonly reported in the literature. However, in practical settings, in particular for atmospheric channels, the quantum channel’s instability complicates the pre-definiton of accepted statistics before the protocol execution, as required by the security argument. This leads to a high abort rate of the protocol and low effective key rates. To address this, in Section IV, we give a variable-length security argument for a variable-length version of our protocol [39], proving again composable finite-size variable-length security against both i.i.d. collective and coherent attacks. Finally, in Section V, we apply our recent method [31] and present both variable-length and fixed-lengh key rates in Section VI.
@@ II Protocol
2.) Measurement. Alice and Bob randomly and independently decide to measure either in their computational bases, represented by POVMs $\{A_{1}^{x}\}_{x=0}^{d-1},\{B_{1}^{y}\}_{y=0}^{d-1}$, or in one of the test bases, represented by POVMs $\{A_{b}^{x}\}_{x=0}^{d-1},\{B_{b}^{y}\}_{y=0}^{d-1}$ for $b\in\{2,...,n_{\mathrm{bases}}\}$, where $n_{\mathrm{bases}}$ denotes the total number of bases used in the protocol, and record their outcomes in their respective private registers $\bar{A}$ and $\bar{B}$. Practically, this can be implemented either passively via beamsplitters, or using active switching with random numbers. We denote the probabilities that basis $i$ is chosen by $p_{i}\geq 0$ and $d$ is the dimension of the protocol.
Steps 1.) and 2.) are repeated $N$ times. Once they finish measuring, Alice may announce a random permutation, which both parties then apply to their data.
@@
4.) Acceptance Testing. The communicating parties disclose a randomly chosen subset consisting of $k_{T}$ of their measurement results over the public channel and perform a so-called acceptance test (see Theorem 1). This test helps them to define a set $\mathcal{S}^{\mathrm{AT}}$. They also may announce to perform subspace postselection.
5.) Key Map Based on their announcements and their measurement results one party performs a key map $g$, which formalizes the process of assigning a logical bit value $\{0,...,d-1\}$ to each of the remaining $N-k_{T}$ rounds as a function of the publicly known announcements and the private measurement data. This step also may allow discarding certain rounds (by assigning the symbol $\perp$ or, more efficiently, by directly shortening the string [39]). In case Alice performs the key map, this is called direct reconciliation; in case it is done by Bob, it is called reverse reconciliation.
@@ III Security Argument
In this work, we aim to demonstrate the composable security of a generic high-dimensional QKD protocol against coherent attacks in the finite-size regime. This task, however, presents several challenges compared to the asymptotic analysis carried out in [31]. While in the asymptotic regime, the law of large numbers justifies assuming that our measurement outcomes are known with certainty, in the finite-size regime we need to thoroughly design and perform an acceptance test. Then, the analysis proceeds on the set of all states that are $\epsilon_{\mathrm{AT}}$-likely to be compatible with our observations. However, the direct use of the measured statistics, as it is common in discrete-variable QKD, for HD-QKD performed in practical measurement setups requires numerical security arguments. As identified in Ref. [27], these methods suffer from serious limitations due to complexity issues, which motivated the development of a new, semi-analytic security argument for practical HD-QKD setups in Ref. [31]. This work avoids the use of full statistics, allowing for direct solutions and bypassing highly demanding numerical convex optimization routines. However, it also means that we have to construct the acceptance set based on suitable coarse-grained statistics. Additionally, we need to take into account that all subroutines used for the QKD protocol may fail with a certain probability, such as error verification failing to align the keys between Alice and Bob or privacy amplification failing to decouple Eve from the final key. Each of these contributions has to be taken into account carefully in a composable security argument. Proving composable finite-size security against coherent attacks is inherently challenging, as most arguments available rely on an assumed i.i.d. structure of the underlying quantum states. Consequently, a lift from collective i.i.d. to coherent attacks is required.
@@
We address these issues as follows. Initially, we consider collective i.i.d. attacks, where Eve prepares a fresh ancilla state that interacts identically with each protocol round. She then may store this ancilla state in her quantum memory and wait until Alice and Bob have finished executing their QKD protocol, including their announcements, before measuring her quantum memory. Mathematically, this implies that different rounds are uncorrelated and that Alice’s and Bob’s shared state has tensor-product structure $\rho=\rho_{AB}^{\otimes N}$. We then adapt an argument by Refs. [40, 41] to our needs and formulate an acceptance test with non-unique acceptance criteria. Based on this test, we prove composable security within Renner’s $\epsilon$-security framework [42]. Finally, we apply the postselection technique [37, 38] to lift our collective i.i.d. key rates to secure key rates against coherent attacks, thereby completing the proof.
@@
The first goal of our security argument is to prove the security of the considered protocol against coherent attacks. We aim to tackle this task in two steps: We first prove security against i.i.d. collective attacks and then we lift our proof to coherent attacks. For our collective attacks argument, we follow the $\epsilon$-security framework by Renner [42]. Thus, for now, let us assume i.i.d. collective attacks, i.e., Alice’s and Bob’s shared quantum state after $N$ rounds of transmission has tensor-product structure $\rho_{AB}=\tau_{AB}^{\otimes N}$. Let $\rho_{ABE}=\tau_{ABE}^{\otimes N}$ be a purification of this state, held by Eve. Alice and Bob then randomly select $k_{T}$ of those rounds for statistical testing. Formally, we now need to construct a set of $\epsilon_{\mathrm{AT}}$-filtered states [42, 40], a set $\left(\mathcal{S}^{\mathrm{AT}}\right)^{\mathsf{c}}$ of quantum states $\sigma$ leading to an accepted statistics with probability less than $\epsilon_{\mathrm{AT}}$, $\mathrm{Pr}\left[\mathrm{accept}|\sigma\right]\leq\epsilon_{\mathrm{AT}}$. In other words, we aim to define a set $\mathcal{S}^{\mathrm{AT}}$ such that all states in $\left(\mathcal{S}^{\mathrm{AT}}\right)^{\mathsf{c}}$ could have generated the accepted statistics only with probability less than $\epsilon_{\mathrm{AT}}$. To define such a set, we adapt the acceptance testing theorem from Ref. [41, Theorem 3].
Theorem 1 (Acceptance Test [41]).
@@
$\mathcal{S}^{\mathrm{AT}}:=\left\{\sigma\in\mathcal{D}(\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{E}):\right.$ (2)
@@
where $r_{X}$ is the $X$th element of the vector $\mathbf{r}$ and likewise for $t_{X}$. For every $X\in\Theta$, let
@@
where $x:=\|X\|_{\infty}$ and $m_{X}$ is the number of tests for the observable $X$. If $\rho\not\in\mathcal{S}^{\mathrm{AT}}$, then the probability of accepting the statistics generated by the i.i.d. measurements of $\rho^{\otimes n}$ is bounded above by $\epsilon_{\mathrm{AT}}$. That is, the complement of $\mathcal{S}^{\mathrm{AT}}$ are all $\epsilon_{\mathrm{AT}}$-filtered.
@@
    $\left(\tau_{AB}\notin\mathcal{S}^{\mathrm{AT}}\right)\land\mathcal{P}_{\mathrm{AT}}$ – the test passes for an input state not being in $\mathcal{S}^{\mathrm{AT}}$,
@@
    $\left(\tau_{AB}\notin\mathcal{S}^{\mathrm{AT}}\right)\land\mathcal{P}_{\mathrm{AT}}^{\mathsf{c}}$ – the test fails for an input state not being in $\mathcal{S}^{\mathrm{AT}}$.
While cases a) and d) correspond to the desired behavior of the test, note that in case of b), although the protocol erroneously aborts, it is nevertheless secure. By design of the acceptance testing theorem, the probability that c) occurs is bounded by $\epsilon_{\mathrm{AT}}$. Additionally, we denote the event that the error-verification routine passes by $\mathcal{P}_{\mathrm{EV}}$ and abbreviate the event of both passing by $\mathcal{P}$.
Composable security builds upon the following Gedankenexperiment [43]. The adversary is given the output of both the real and an idealized version of the QKD protocol and aims to distinguish between both protocols without having access to Alice’s and Bob’s private data. The ideal output is uniformly distributed and completely decoupled from Eve, $\rho_{\mathrm{ideal}}:=\pi_{K_{A}K_{B}}\otimes\rho_{E}$, where $\pi_{K_{A}K_{B}}:=\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}\outerproduct{k}{k}\otimes\outerproduct{k}{k}$ |. The output of the real protocol shall be denoted by $\rho_{K_{A}K_{B}E}:=\Phi(\rho_{ABE})$, where $\Phi$ denotes the map representing the action of the protocol (which will be detailed on in Section V.2). Then, we aim to bound the distinguishing advantage of an adversary,
$\frac{1}{2}\left|\left|\rho_{K_{A}K_{B}E}-\pi_{K_{A}K_{B}}\otimes\rho_{E}\right|\right|_{1}\leq\epsilon.$ (3)
@@
$\frac{1}{2}\left|\left|\rho_{K_{A}K_{B}E}-\pi_{K_{A}K_{B}}\otimes\rho_{E}\right|\right|_{1}$
$=$ $0+\frac{1}{2}\left|\left|\rho_{K_{A}K_{B}E\land\mathcal{P}}-\pi_{K_{A}K_{B}}\otimes\rho_{E\land\mathcal{P}}\right|\right|_{1}$
$\leq$ $\frac{1}{2}\left|\left|\rho_{K_{A}K_{B}E\land\mathcal{P}}-\rho_{K_{A}K_{A}E\land\mathcal{P}}\right|\right|_{1}$
$+\frac{1}{2}\left|\left|\rho_{K_{A}K_{A}E\land\mathcal{P}}-\pi_{K_{A}K_{B}}\otimes\rho_{E\land\mathcal{P}}\right|\right|_{1}$
$\leq$ $\epsilon_{\mathrm{EV}}+\frac{1}{2}\left|\left|\rho_{K_{A}K_{A}E\land\mathcal{P}}-\pi_{K_{A}K_{B}}\otimes\rho_{E\land\mathcal{P}}\right|\right|_{1},$
where we used $\mathrm{Pr}\left[\mathcal{P}\right]\sigma_{A|\mathcal{P}}=:\sigma_{A\land\mathcal{P}}$. For the first inequality, we applied the triangle inequality and for the second inequality we used that the first term is bounded by $\epsilon_{\mathrm{EV}}$. To ease notation, we now drop Bob’s key register, as it is equal to Alice’s. For the second term, we use the fact that the remaining cases a) and c) are mutually exclusive, which leads to the upper bound
$\frac{1}{2}\left|\left|\rho_{K_{A}K_{B}E\land\mathcal{P}}-\pi_{K_{A}K_{B}}\otimes\rho_{E\land\mathcal{P}}\right|\right|_{1}$
@@
$\mathrm{Pr}\left[\tau_{AB}\notin\mathcal{S}^{\mathrm{AT}}\right]\frac{1}{2}\left|\left|\rho_{K_{A}E\land\mathcal{P}}-\pi_{K_{A}}\otimes\rho_{E\land\mathcal{P}}\right|\right|_{1}$
@@
While we now have successfully treated cases b) - d), we are left with case a), where Alice’s and Bob’s shared key is not fully private with respect to Eve. To decouple their key from Eve, the protocol requires Alice and Bob to perform privacy amplification. The leftover hashing lemma (LHL) [42, Corollary 5.6.1] relates the length of the hashed key to the security parameter,
$\frac{1}{2}\left|\left|\rho_{K_{A}E\land\mathcal{P}}-\pi_{K_{A}}\otimes\rho_{E\land\mathcal{P}}\right|\right|_{1}$
@@
where $C$ denotes the transcript of the information reconciliation phase. Thus, as long as the length of the hashed key satisfies
$\ell\leq H_{\mathrm{min}}^{\tilde{\epsilon}}(X|EC)_{\Phi^{(n)}(\rho_{ABE})\land\mathcal{P}}+2\log_{2}\left(\epsilon_{\mathrm{sec}}-2\tilde{\epsilon}\right),$
the key is secret with secrecy parameter $\epsilon_{\mathrm{sec}}$. Next, we use [44, Lemma 10] $H_{\mathrm{min}}^{\tilde{\epsilon}}(X|EC)_{\sigma\land\mathcal{P}}\geq H_{\mathrm{min}}^{\tilde{\epsilon}}(X|EC)_{\sigma}$ as well as Lemma 6.4.1 in [42] to remove the information reconciliation transcript from the smooth min-entropy at the cost of introducing a leakage term $\mathrm{leak}_{\mathrm{IR}}$ and arrive at
$\ell\leq H_{\mathrm{min}}^{\tilde{\epsilon}}(X|E)_{\Phi^{(n)}(\rho_{ABE})}-\mathrm{leak}_{\mathrm{IR}}+2\log_{2}\left(\epsilon_{\mathrm{sec}}-2\tilde{\epsilon}\right).$
@@
$+2\log_{2}\left(\epsilon_{\mathrm{sec}}-2\tilde{\epsilon}\right).$
Let $n:=N-k_{T}$ denote the number of rounds used for key generation. Since the smooth min-entropy of a quantum state is defined as the maximum of the min-entropy over all quantum states in an $\epsilon$-ball around the considered quantum state, we can lower bound $H_{\mathrm{min}}^{\tilde{\epsilon}}(X|E)_{\Phi(\rho_{ABE})}\geq H_{\mathrm{min}}(X|E)_{\Phi(\rho_{ABE})}$ (i.e., set $\tilde{\epsilon}=0$). Next, we use the fact that the min-entropy is additive for i.i.d. copies [45, Table 4.1] and obtain
@@
Thus, we choose $\epsilon_{\mathrm{sec}}=\epsilon_{\mathrm{PA}}$ and obtain a valid lower bound on the secure key rate as long as we choose $\ell$ such that

### v2

@@ Abstract
Practical implementations of Quantum Key Distribution (QKD) extending beyond urban areas commonly use satellite links. However, the transmission of quantum states through the Earth’s atmosphere is highly susceptible to noise, restricting its application primarily to nighttime. High-dimensional (HD) QKD offers a promising solution to this limitation by employing high-dimensionally entangled quantum states. Although experimental platforms for HD QKD exist, previous security analyses have been limited to the asymptotic regime and have either relied on impractical measurements or employed computationally demanding convex optimization tasks restricting the security analysis to low dimensions. In this work, we bridge this gap by presenting a composable finite-size security proof against both collective and coherent attacks for a general HD QKD protocol that uses only experimentally accessible measurements. In addition to the conventional, yet impractical ‘one-shot’ key rates, we provide a practical variable-length security argument that yields significantly higher expected key rates. This approach is particularly crucial under rapidly changing and turbulent atmospheric conditions, as encountered in free-space and satellite-based QKD platforms.
@@ I Introduction
Quantum Key Distribution [Bennett_Brassard_1984, Ekert_1991] enables two remote parties, commonly called Alice and Bob, to establish cryptographically secure keys, without relying on assumptions on the hardness of mathematical problems. Over recent decades, significant advancements in both experimental techniques and security proofs have evolved QKD from an interesting scientific concept toward one of the most developed quantum technologies [Scarani_2009, Pirandola_2020, Usenko_2025], and it is now on the edge of practical use for both private and governmental applications. Implementations range from fiber links [Hiskett_2006, SECOQC_2009, Rosenberg_2007, Peng_2007, Boaron_2018] for urban areas to free-space [Manderbach_2007, Krzic_2023] and satellite-based links [Qubesat_2018, Micius_2022, QEYSSAT_2023, Sidhu_2023, Orsucci_2024] that span countries and even continents. Although the transmission of quantum signals through the atmosphere offers a more favorable loss profile compared to fiber-based systems, it also suffers from high noise sensitivity, which essentially restricts operation to nighttime hours.
@@
High-dimensional quantum states [Kwiat_1997, Barreiro_2005, Martin_2017] have proven to enhance noise tolerance in entanglement distribution tasks [Ecker_2019], making high-dimensional quantum key distribution (HD QKD) a promising candidate for improving noise robustness in cryptographic tasks. Nonetheless, the security analysis of practically implementable HD QKD protocols is so far rather uncharted. Previous works have primarily focused on asymptotic security, employing either analytical methods [Sheridan_2010, Ferenczi_2012, Doda_2021, Wyderka_2025, Thomas_2025] that leverage symmetry arguments or numerical approaches [BKPH_2023] involving convex optimization [Coles_2016, Winick_2018, Araujo_2022]. Analytical arguments often require $2$ or even $d+1$ mutually unbiased bases (MUB) measurements, which are impractical, particularly for large dimensions. Numerical methods avoid this but require solving very high-dimensional convex optimization problems, which are computationally demanding and very memory-intensive, rendering them infeasible even for low to medium dimensions [BKPH_2023]. Our recent work in Ref. [Kanitschar_2024a] has closed this gap between experimentally available setups [Bavaresco_2018, Hu_2020, Schneeloch_2019, Ponce_2022, Maxwell_2022] and asymptotic security, by providing an efficient semianalytic method to evaluate key rates for practical HD QKD setups, relying solely on practically implementable measurements. Yet, the security of high-dimensional QKD protocols in the finite-size regime remains an open question.
@@
In this study, we address this issue by presenting a full, composable finite-size security proof for general high-dimensional QKD protocols. We begin by outlining the general HD QKD protocol in Section II. Section III details our statistical testing procedure and establishes composable finite-size security against i.i.d. collective attacks. We then apply the postselection technique [Christandl_2009, Nahar_2024] to extend security to coherent attacks. These secure key rates refer to the so-called fixed-length scenario, which is commonly reported in the literature. However, in practical settings, in particular for atmospheric channels, the quantum channel’s instability complicates the pre-definition of accepted statistics before the execution of the protocol, as required by the security argument. This leads to a high abort rate of the protocol and low effective key rates. To address this, in Section IV, we give a variable-length security argument for a variable-length version of our protocol [Tupkary_2024], again proving composable finite-size variable-length security against both i.i.d. collective and coherent attacks. Finally, in Section V, we apply our recent method [Kanitschar_2024a] and present both variable-length and fixed-length key rates in Section VI.
@@ II Protocol
2.) Measurement. Alice and Bob randomly and independently decide to measure either in their computational bases, represented by POVMs $\{A_{1}^{x}\}_{x=0}^{d-1},\{B_{1}^{y}\}_{y=0}^{d-1}$, or in one of the test bases, represented by POVMs $\{A_{b}^{x}\}_{x=0}^{d-1},\{B_{b}^{y}\}_{y=0}^{d-1}$ for $b\in\{2,...,n_{\mathrm{bases}}\}$, where $n_{\mathrm{bases}}$ denotes the total number of bases used in the protocol, and record their outcomes in their respective private registers $\bar{A}$ and $\bar{B}$. Practically, this can be implemented either passively via beamsplitters or using active switching with random numbers. We denote the probability that basis $i$ is chosen by $p_{i}\geq 0$, and $d$ is the dimension of the protocol.
Steps 1.) and 2.) are repeated $N$ times. Once the measurements are completed, Alice may announce a random permutation, which both parties then apply to their data.
@@
4.) Acceptance Testing. The communicating parties disclose a randomly chosen subset consisting of $k_{T}$ of their measurement results over the public channel and perform a so-called acceptance test (see Theorem 1). This test helps them to define a set $\mathcal{S}^{\mathrm{AT}}$. They may also announce that they will perform subspace postselection.
5.) Key Map Based on their announcements and their measurement results one party performs a key map $g$, which formalizes the process of assigning a logical bit value $\{0,...,d-1\}$ to each of the remaining $N-k_{T}$ rounds as a function of the publicly known announcements and the private measurement data. This step also may allow discarding certain rounds (by assigning the symbol $\perp$ or, more efficiently, by directly shortening the string [Tupkary_2024]). In case Alice performs the key map, this is called direct reconciliation; in case it is done by Bob, it is called reverse reconciliation.
@@ III Security Argument
In this work, we aim to demonstrate the composable security of a generic high-dimensional QKD protocol against coherent attacks in the finite-size regime. This task, however, presents several challenges compared to the asymptotic analysis carried out in Ref. [Kanitschar_2024a]. Although in the asymptotic regime, the law of large numbers justifies assuming that our measurement outcomes are known with certainty, in the finite-size regime, we need to thoroughly design and perform an acceptance test. The acceptance test carefully compares the observations with the expected channel behaviour. If the test passes, the analysis proceeds on the set of all states that are $\epsilon_{\mathrm{AT}}$-likely to be compatible with our observations. Instead of considering a plethora of attack strategies that Eve could follow (a nearly impossible task), this allows us to base the security argument solely on the observations made by the communicating parties. However, the direct use of the measured statistics, as it is common in discrete-variable QKD, for HD-QKD performed in practical measurement setups requires numerical security arguments. As identified in Ref. [BKPH_2023], these methods suffer from serious limitations due to complexity issues, which motivated the development of a new, semi-analytic security argument for practical HD-QKD setups in Ref. [Kanitschar_2024a]. This work avoids the use of full statistics, allowing for direct solutions and bypassing highly demanding numerical convex optimization routines. However, it also means that we have to construct the acceptance set based on suitable coarse-grained statistics. Additionally, we need to take into account that all subroutines used for the QKD protocol may fail with a certain probability, such as error verification failing to align the keys between Alice and Bob or privacy amplification failing to decouple Eve from the final key. Each of these contributions has to be taken into account carefully in a composable security argument. Proving composable finite-size security against coherent attacks is inherently challenging, as most available arguments rely on an assumed i.i.d. structure of the underlying quantum states. Consequently, a lift from collective i.i.d. to coherent attacks is required.
@@
We address these issues as follows. Initially, we consider collective i.i.d. attacks, where Eve prepares a fresh ancilla state that interacts identically with each protocol round. She then may store this ancilla state in her quantum memory and wait until Alice and Bob have finished executing their QKD protocol, including their announcements, before measuring her quantum memory. Mathematically, this implies that different rounds are uncorrelated and that Alice’s and Bob’s shared state has tensor-product structure $\rho=\rho_{AB}^{\otimes N}$. We then adapt an argument by Refs. [George_2021, Kanitschar_2023] to our needs and formulate an acceptance test with non-unique acceptance criteria. Based on this test, we prove composable security within Renner’s $\epsilon$-security framework [Renner_2005]. Finally, we apply the postselection technique [Christandl_2009, Nahar_2024] to lift our collective i.i.d. key rates to secure key rates against coherent attacks, thereby completing the proof.
@@
The first goal of our security argument is to prove the security of the considered protocol against coherent attacks. We aim to tackle this task in two steps: We first prove security against i.i.d. collective attacks, and we then lift our proof to coherent attacks. For our collective attacks argument, we follow the $\epsilon$-security framework by Renner [Renner_2005]. Thus, for now, let us assume i.i.d. collective attacks, i.e., Alice’s and Bob’s shared quantum state after $N$ rounds of transmission has tensor-product structure $\rho_{AB}=\tau_{AB}^{\otimes N}$. Let $\rho_{ABE}=\tau_{ABE}^{\otimes N}$ be a purification of this state, held by Eve. Alice and Bob then randomly select $k_{T}$ of those rounds for statistical testing.
Formally, this means we need to construct a set of $\epsilon_{\mathrm{AT}}$-filtered states [Renner_2005, George_2021], which is a set $\left(\mathcal{S}^{\mathrm{AT}}\right)^{\mathsf{c}}$ of quantum states $\sigma$ leading to an accepted statistics with probability less than $\epsilon_{\mathrm{AT}}$, $\mathrm{Pr}\left[\mathrm{accept}|\sigma\right]\leq\epsilon_{\mathrm{AT}}$. In other words, our goal is to define a set $\mathcal{S}^{\mathrm{AT}}$ such that all states in its complement $\left(\mathcal{S}^{\mathrm{AT}}\right)^{\mathsf{c}}$ could have generated the accepted statistics only with probability less than $\epsilon_{\mathrm{AT}}$. Intuitively, this allows us to split the set of all quantum states into two subsets: the set of states that are very unlikely to have produced the observed statistics, and the set of states that have potentially caused our observations. We formalise this notion in the following theorem, which is an adapted version of Ref. [Kanitschar_2023, Theorem 3].
Theorem 1 (Acceptance Test [Kanitschar_2023]).
@@
$\mathcal{S}^{\mathrm{AT}}=\left\{\sigma\in\mathcal{D}(\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{E}):\right$. (2)
@@
where $r_{X}$ is the $X$-th element of the vector $\mathbf{r}$ and likewise for $t_{X}$. For every $X\in\Theta$, let
@@
where $x:=\|X\|_{\infty}$ and $m_{X}$ is the number of tests for the observable $X$. If $\rho\not\in\mathcal{S}^{\mathrm{AT}}$, then the probability of accepting the statistics generated by the i.i.d. measurements of $\rho^{\otimes n}$ is bounded above by $\epsilon_{\mathrm{AT}}$. That is, the complement of $\mathcal{S}^{\mathrm{AT}}$ consists of all $\epsilon_{\mathrm{AT}}$-filtered states.
@@
    $\left(\tau_{AB}\notin\mathcal{S}^{\mathrm{AT}}\right)\land\mathcal{P}_{\mathrm{AT}}$ – the test passes for an input state not in $\mathcal{S}^{\mathrm{AT}}$,
@@
    $\left(\tau_{AB}\notin\mathcal{S}^{\mathrm{AT}}\right)\land\mathcal{P}_{\mathrm{AT}}^{\mathsf{c}}$ – the test fails for an input state not in $\mathcal{S}^{\mathrm{AT}}$.
While cases a) and d) correspond to the desired behavior of the test, note that in case b), although the protocol erroneously aborts, it is nevertheless secure. By design of the acceptance testing theorem, the probability that c) occurs is bounded by $\epsilon_{\mathrm{AT}}$. Additionally, we denote the event that the error-verification routine passes by $\mathcal{P}_{\mathrm{EV}}$ and abbreviate the event of both passing by $\mathcal{P}$.
Composable security builds upon the following Gedankenexperiment [Portmann_2022]. The adversary is given the output of both the real and an idealized version of the QKD protocol and aims to distinguish between both protocols without having access to Alice’s and Bob’s private data. The ideal output is uniformly distributed and completely decoupled from Eve, $\rho_{\mathrm{ideal}}:=\pi_{K_{A}K_{B}}\otimes\rho_{E}$, where $|kk|\pi_{K_{A}K_{B}}:=\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}\outerproduct$$\otimes\outerproduct{k}{k}$. The output of the real protocol shall be denoted by $\rho_{K_{A}K_{B}E}:=\Phi(\rho_{ABE})$, where $\Phi$ denotes the map representing the action of the protocol (which will be detailed on in Section V.2). Then, we aim to bound the distinguishing advantage of an adversary,
[…]
@@
[…]
where we used $\mathrm{Pr}\left[\mathcal{P}\right]\sigma_{A|\mathcal{P}}=:\sigma_{A\land\mathcal{P}}$. For the first inequality, we applied the triangle inequality, and for the second inequality, we used that the first term is bounded by $\epsilon_{\mathrm{EV}}$. To ease notation, we now drop Bob’s key register, as it is equal to Alice’s. For the second term, we use the fact that the remaining cases a) and c) are mutually exclusive, which leads to the upper bound
[…]
@@
[…]
@@
While we have now successfully treated cases b) - d), we are left with case a), where Alice’s and Bob’s shared key is not fully private with respect to Eve. To decouple their key from Eve, the protocol requires Alice and Bob to perform privacy amplification. The leftover hashing lemma (LHL) [Renner_2005, Corollary 5.6.1] relates the length of the hashed key to the security parameter,
[…]
@@
where $C$ denotes the transcript of the information reconciliation phase. Thus, provided that the length of the hashed key satisfies
[…]
the key is secret with secrecy parameter $\epsilon_{\mathrm{sec}}$. Next, we use [Tomamichel_2017, Lemma 10] $H_{\mathrm{min}}^{\tilde{\epsilon}}(X|EC)_{\sigma\land\mathcal{P}}\geq H_{\mathrm{min}}^{\tilde{\epsilon}}(X|EC)_{\sigma}$ as well as Lemma 6.4.1 in [Renner_2005] to remove the information reconciliation transcript from the smooth min-entropy at the cost of introducing a leakage term $\mathrm{leak}_{\mathrm{IR}}$ and arrive at
[…]
@@
[…]
Let $n:=N-k_{T}$ denote the number of rounds used for key generation. Since the smooth min-entropy of a quantum state is defined as the maximum of the min-entropy over all quantum states in an $\epsilon$-ball around the considered quantum state, we can lower bound $H_{\mathrm{min}}^{\tilde{\epsilon}}(X|E)_{\Phi(\rho_{ABE})}\geq H_{\mathrm{min}}(X|E)_{\Phi(\rho_{ABE})}$ (i.e., set $\tilde{\epsilon}=0$). Next, we use the fact that the min-entropy is additive for i.i.d. copies [Thomamichel_Thesis, Table 4.1] and obtain
@@
Thus, we choose $\epsilon_{\mathrm{sec}}=\epsilon_{\mathrm{PA}}$ and obtain a valid lower bound on the secure key rate provided that we choose $\ell$ such that
@@
Leaving aside the technical details, for pre-fixed subroutine security parameters $\epsilon_{\mathrm{EV}},\epsilon_{\mathrm{PA}}$ and $\epsilon_{\mathrm{AT}}$ this theorem relates the achievable secure key length $\ell$ under collective attacks to an optimisation problem over the set $\mathcal{S}^{\mathrm{AT}}$, which we defined earlier in the acceptance test, along with correction terms due to error correction and privacy amplification. It turns out that the remaining optimisation problem can now be solved using the method we introduced in our companion paper [Kanitschar_2024a].